OpenStudy (anonymous):

How to integrate 1/(3-x)^(1/2)

hartnn (hartnn):

@Kuroten87 Hi , \(\Huge \mathcal{\text{Welcome To OpenStudy}\ddot\smile} \) Did you try substitution u= 3-x ?

OpenStudy (anonymous):

I actually tried subbing the square root of 3-x and finding x, then subbing that

OpenStudy (anonymous):

So x would equal 3-u^2 and dx = -2udu

hartnn (hartnn):

ok, now try, u= 3-x (instead of sqrt (3-x))

OpenStudy (anonymous):

Now I have negative integral of 1/(u^1/2)du

hartnn (hartnn):

thats quite simple right ? can you do it ?

OpenStudy (anonymous):

Maybe if I stare at it long enough. Thanks:)

hartnn (hartnn):

lol......let me give you a hint : \(\large \dfrac{1}{x^n}=x^{-n}\) and you have standard integration formula for x^n dx

hartnn (hartnn):

and welcome ^_^

OpenStudy (anonymous):

Lol oh yeah... Basic stuff. My bad.

hartnn (hartnn):

no problem :)

hartnn (hartnn):

oh, and since you are new here, if you have any doubts browsing this site, you can ask me, or to an ambassador (with an encircled A before their name). Have fun learning with Open Study! :)

OpenStudy (anonymous):

Cool! Thanks again!